OpenStudy (anonymous):

metal!!!!! What is the linear velocity of a wheel with diameter of 7.0 m rolling at a rate of 6.0 π rad/sec? a) 66 m/s b) 21 m/s c) 5.4 m/s d) 19 m/s e) 130 m/s

OpenStudy (anonymous):

Ello mate. You need to work out how many revolutions per second the wheel is traveling at, and what distance one revolution is. You can then multiply these two values together to find the velocity in meters per second. For the revolutions per second, we have to know that one revolution is 2pi radians. So 6 pi radians per second must be 3 revolutions per second. For the distance per revolution, we use the circumference of the wheel. wheel circumference = pi*d = pi*7 = 21.99 m 21.99 * 3 = 65.9 m/s, or approximately 66 m/s, so the answer is a)